An ac voltmeter, which displays the rms voltage between the two points touched by its leads, is used to measure voltages in the circuit. In this circuit, the ac generator has an rms voltage of 8.90 V and a frequency of 25.0 kHz . The inductance in the circuit is 0.250 mH , the capacitance is 0.150 μF, and the resistance is 4.60 Ω.

Part A:

What is the reading on a voltmeter when it is connected to points A and B?

Express your answer using two significant figures.

Part B:

What is the reading on a voltmeter when it is connected to points B and C?

Express your answer using two significant figures.

Part C:

What is the reading on a voltmeter when it is connected to points A and C?

Express your answer using two significant figures.

Part D:

What is the reading on a voltmeter when it is connected to points A and D?

Express your answer using two significant figures.
Answer #1

Solution in the uploaded images

frast XL = 1.58 x 39.25 Solution. L R A B C mmit Vans 8.00u V = 25 kHz L= 0.250mt C = 0.150 MF R=4.6or Wa 278 = 21 25x103 = 1Part B: Ugc = Ipas x Xe = 1.58*42.46= 67.51 68u KERS Port Ci VAC = Ipas (XL+X c) = 1:59 (39.25 + 42.46) - 129.92 v ~ 13ou Par